# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

### v1

@@ 4.1 Repetition Line Experiment
The results in Table 12 reveal a consistent trend of marked improvements in nearly all models with the improvements ranging from 1% to an upmost of 14%. Interestingly, the models that had the largest improvements (>10%) are all Mistral-based, which suggests that this particular family of models, whether for training reasons or how they were initially trained, seem to prefer reiterating the instructions or the important information in a prompt. Other model families generally still had respectable improvements of around 1-2%, which suggests that even other models prefer reiterating certain parts of the prompt, with only Claude 3 Sonnet and GPT-3.5-Turbo having a decrease in performance.
@@ 4.2 Knowledge Reinforcement Experiment
In this experiment, we sought to determine whether providing models with information they had previously generated would enhance their performance. This idea stemmed from an earlier study (not included in this paper) where we provided models with the DSM-5 criteria for depression. Interestingly, some models improved in accuracy, while others experienced a decline. To investigate this phenomenon more systematically, we designed the Knowledge Reinforcement Experiment.
The experiment was conducted in two steps. First, each LLM was prompted to list symptoms it associates with depression. The prompt used for this step is presented in Table 13.
Table 13: Prompt used for extracting disorder symptoms from LLM s
Give me a detailed list of symptoms that indicate whether a person is depressed or not.
@@
Table 15 presents the results of this experiment. The main goal is to observe whether a model would benefit from being provided information it very clearly knows, given that it provided us that information to begin with.
The results lack a clear pattern, with some models showing a moderate increase in accuracy, notably GPT-3.5-Turbo and Claude 3 Sonnet. However, most smaller LLM s experienced a decrease in accuracy, some quite significantly. We are unsure why smaller LLM s were negatively affected, as the additional information should have been neutral or even beneficial. One hypothesis is that the lengthy prompts with extensive information might have distracted them from the main instructions and task, potentially leading to poorer performance.
Table 15: Results of the Knowledge Reinforcement experiment compared to the baseline BIN-1 prompt results
@@
Gemini 1.0 Pro 71.4% +0.2%
@@
Mixtral 8x22b 71.9% -3.2%
Llama 3 70b 71.9% +0.9%
Llama 2 70b 75% -4.1%
Mixtral 8x7b 69.3% -4.3%
Llama 3 8b 66.8% -8.5%
Gemma 7b 55% 0%
Mistral 7b 64.4% -0.6%
5 Challenges, Conclusions and Future Directions
5.1 Challenges
While the potential of LLM s in the realm of mental health is promising, our investigation into their capabilities for binary disorder detection faced several challenges. These obstacles arose from both the limitations of available datasets and the inherent characteristics of the LLM s themselves. Understanding these challenges is crucial for interpreting the results of our evaluation and guiding future research in this domain.
5.1.1 Datasets Based Challenges
The ideal evaluation would involve clinical reports containing patient summaries and diagnoses, as these most accurately reflect an individual’s physical and mental health. However, due to the confidential nature of such data, access to these datasets is either extremely difficult or nearly impossible without a close working relationship with a hospital or mental health institution. As a result, we relied on publicly available social media datasets, which, while more accessible, are inherently less accurate.
These social media datasets presented quality concerns. Many labels were based on context (e.g., subreddit of origin), leading to potential noise in the data. For instance, posts from r/Depression may not always indicate depression, but could be supportive messages or questions. We prioritized datasets with human supervision to mitigate this issue.
However, even with human-labeled data, psychiatric evaluation remains a subjective process. Trained professionals may have differing opinions, which can introduce variability into the data. As a result, even datasets labeled by experts may still contain some degree of noise due to these differences in interpretation.
5.1.2 LLM Based Challenges
The use of LLM s presents several challenges. One significant issue is the cost associated with models like GPT-4 and Claude Opus, which limited our experiments to samples from test sets. This high expense raises concerns about the scalability of these models for real-world applications.
Another challenge is the context limit of some models, such as Phi-2, which has a 2048-token limit. This constraint made it difficult to handle datasets with very long posts. However, with newer models featuring larger context windows, this limitation is expected to become less of a concern in the future.
API filtering presents a significant challenge. For example, Gemini 1.0 Pro’s API has stringent filtering policies that block prompts mentioning sensitive topics such as pedophilia, sexual abuse, or grooming. This restriction limits the model’s applicability in mental health contexts, where discussing sensitive issues is often necessary. Consequently, this filtering issue was the primary reason other Gemini models, such as Gemini 1.5 Flash/Pro, were not tested in our experiments.
Ethical guards implemented in some models, such as the Llama 3.1, have also been problematic. These safeguards can significantly hinder performance in mental health contexts compared to their older Llama 3 counterparts.
Additionally, some models exhibit stubborn behavior, resisting the provision of explanations despite clear instructions. Mistral models, in particular, have shown a higher level of stubbornness compared to GPT models. This inconsistency in following instructions is concerning and may be linked to the training data or processes used.
Finally, the sensitivity of models to prompt changes has been another issue. Even minor modifications to prompts can lead to excessive sensitivity. For example, adding redundant information notably affected the performance of models like Llama 3 70b and Gemini 1.0 Pro, highlighting the importance of careful prompt design.
@@ 5.2 Future Directions
Building upon the insights gained from our investigation into the use of LLM s for mental health applications, several promising avenues for future research have emerged. These directions are aimed at enhancing the interpretability, performance, and applicability of LLM s in real-world mental health contexts, while also addressing some of the key challenges identified in our study.
One promising area of exploration is the application of CoT prompting to enhance the interpretability and decision-making capabilities of LLM s in psychiatric contexts. CoT prompting involves structuring model prompts in a way that encourages step-by-step reasoning, which could be particularly beneficial for understanding complex mental health issues. This approach could help clarify the rationale behind model outputs, making it easier to validate and trust the decisions made by LLM s. Investigating how CoT prompting affects the accuracy and reliability of mental health diagnostics will be crucial for deploying LLM s in real-world psychiatric applications.
Another innovative approach to improving LLM performance involves the use of Retrieval-Augmented Generation (RAG) techniques. By integrating LLM s with a database of relevant knowledge, such as books, case studies, or even anonymized medical records, RAG could provide additional context and grounding to the model’s responses. This method can enhance the model’s ability to generate more informed and contextually accurate outputs, thus improving diagnostic accuracy and relevance.
Fine-tuning LLM s on specific psychiatric datasets represents another important area for future investigation. Tailoring models to domain-specific data could potentially improve their performance and generalizability across different mental health conditions. Evaluating the impact of fine-tuning on diagnostic accuracy and model robustness will provide insights into how LLM s can be optimized for diverse psychiatric tasks and populations.
Expanding the evaluation of LLM s to include other psychiatric tasks, such as dialogue safety and empathy, is essential for a comprehensive understanding of their capabilities. Publicly available datasets for these tasks can provide valuable insights into how well LLM s perform beyond binary disorder detection. Additionally, examining how LLM s handle the diagnosis of multiple co-occurring mental disorders could provide a more nuanced view of their diagnostic capabilities and limitations.
Furthermore, future research should include evaluating LLM s with datasets in other languages, as well as those involving different modalities such as audio, images, and videos. With advancements in multimodal models like GPT-4o and Llama 3.1, integrating and analyzing data from diverse sources can enhance diagnostic accuracy and offer a richer understanding of mental health conditions across various contexts.
By pursuing these directions, researchers can address current challenges and push the boundaries of LLM applications in mental health, ultimately leading to more effective and ethically sound tools for psychiatric care.
5.3 Conclusions
In this study, we conducted an extensive evaluation of the capabilities of LLM s in the context of mental health tasks, specifically focusing on the detection and evaluation of mental disorders from social media posts. Our investigation encompassed a wide range of LLM s, including state-of-the-art models like GPT-4 and Claude 3.5, as well as smaller, more accessible models such as Gemma 2, Phi-3, LLama 3/3.1 and others. We employed diverse prompt templates and evaluation metrics to assess their performance across three primary tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment.
Our findings revealed that LLM s hold significant promise in assisting with mental health tasks. In the binary disorder detection task, models like GPT-4 and Llama 3 exhibited superior performance, achieving accuracies of up to 85% on certain datasets. The choice of prompt template played a crucial role, with some models favoring more structured prompts while others performed better with open-ended ones. The disorder severity evaluation task highlighted the benefits of few-shot learning, where providing models with a few examples before making predictions led to improved performance in most cases. In the psychiatric knowledge assessment task, recent models generally outperformed older, larger counterparts, demonstrating the advancements in model architecture and training data.
However, the study also identified challenges that must be addressed to effectively deploy LLM s in mental health applications. The limitations of social media datasets, such as potential noise and subjectivity in labeling, and the high costs associated with large models, pose significant concerns. Additionally, ethical restrictions imposed by LLM providers on sensitive queries complicate accurate evaluations in mental health contexts. Despite these hurdles, the potential of LLM s in mental healthcare is undeniable. Future research should focus on improving interpretability and decision-making through techniques like chain-of-thought prompting and retrieval-augmented generation. Additionally, fine-tuning models on specific psychiatric datasets and exploring multimodal data integration can further enhance their utility in this critical domain.
6 Acknowledgement
We gratefully acknowledge and thank the authors of [29] for providing the DEPTWEET dataset.

### v2

@@ 6.1 Repetition Line Experiment
The results in Table 13 reveal a consistent trend of marked improvements in nearly all models with the improvements ranging from 1% to an utmost of 14%. Interestingly, the models that had the largest improvements (>10%) are all Mistral-based, which suggests that this particular family of models, whether for training reasons or how they were initially trained, seem to prefer reiterating the instructions or the important information in a prompt. Other model families generally still had respectable improvements of around 1-2%, which suggests that even other models prefer reiterating certain parts of the prompt, with only Claude 3 Sonnet and GPT-3.5-Turbo having a decrease in performance.
@@ 6.2 Knowledge Reinforcement Experiment
In this experiment, we sought to determine whether providing models with information they had previously generated would enhance their performance. This idea stemmed from an earlier study (not included in this paper) where we provided models with the DSM-Vcriteria for depression. Interestingly, some models improved in accuracy, while others experienced a decline. To investigate this phenomenon more systematically, we designed the Knowledge Reinforcement Experiment.
The experiment was conducted in two steps. First, each LLMwas prompted to list symptoms it associates with depression. In this step, we used "Give me a detailed list of symptoms that indicate whether a person is depressed or not." as a prompt.
@@
Table 15 presents the results of this experiment. The main goal is to observe whether a model would benefit from being provided information it very clearly knows, given that it provided us that information to begin with. The results lack a clear pattern, with some models showing a moderate increase in accuracy, notably GPT-3.5-Turbo and Claude 3 Sonnet. However, most smaller LLMs experienced a decrease in accuracy, some quite significantly. We are unsure why smaller LLMs were negatively affected, as the additional information should have been neutral or even beneficial. One hypothesis is that the lengthy prompts with extensive information might have distracted them from the main instructions and task, potentially leading to poorer performance.
Table 15: Results of the Knowledge Reinforcement experiment compared to the baseline BIN-1.1 prompt on DepSeverity (Binary - 1000 Samples)
@@
Llama 3 8b 66.8% -8.5%
Llama 2 70b 75% -4.1%
Llama 3 70b 71.9% +0.9%
Mistral 7b 64.4% -0.6%
Mixtral 8x7b 69.3% -4.3%
Mixtral 8x22b 71.9% -3.2%
@@
6.3 Clinical Interpretation of Model Performance
While our evaluation primarily reports quantitative performance metrics, the patterns observed also yield clinically relevant insights into why certain models outperform others. These factors extend beyond raw accuracy and should inform model selection for psychiatric applications. Models such as GPT-4o and GPT-4 consistently demonstrated both higher adherence to task instructions and superior accuracy across multiple datasets, particularly in structured classification tasks (e.g., 85.0% accuracy on DEPTWEET using BIN-4). This suggests that strong instruction-following finetuning an approach that explicitly optimizes for compliance with structured prompts may align more closely with the rigid formats used in psychiatric assessments and screening tools. In a clinical workflow, this could translate to more reliable completion of standardized diagnostic questionnaires or risk assessments without unintended deviations. Our results challenge the assumption that larger models automatically yield better clinical performance. For example, the Phi-3-mini (3.8B) model outperformed its larger sibling, Phi-3-small (7B), on the RED SAM dataset (57.8% vs. 57.0%), and Gemma 2 9B outperformed Gemma 2 27B on 5 out of 6 datasets, including an advantage on DEPTWEET (80.4% vs. 77.9%). These findings highlight that architecture, optimization, and fine-tuning objectives can outweigh parameter count, especially when tasks require precise adherence to specialized prompts rather than broad world knowledge. Although specific training datasets for proprietary models are undisclosed, performance patterns suggest that exposure to medical or scientific text either explicitly or through general web-scale sources may enhance psychiatric task performance. Models with likely greater medical content coverage, such as GPT-4o and Qwen 2 72B, showed consistently strong results in psychiatric knowledge assessments (e.g., 91.2% on MedMCQA for Llama 3.1 405B, which is likely influenced by targeted domain adaptation). This reinforces the value of domain-specific fine-tuning or continual pretraining when deploying LLMs for clinical reasoning tasks. We observed that certain models, such as Llama 3.1 and Gemini, refused to process a notable portion of psychiatric task prompts, likely due to strengthened safety alignment. While such refusals limit benchmarking completeness, they may be desirable in real-world triage contexts especially for high-risk content like active suicidality where containment and safe handoff to a human clinician are preferable to automated classification. Clinically, the optimal balance between responsiveness and safety will depend on deployment context: high-throughput population screening may tolerate fewer refusals, whereas acute care settings may benefit from stricter safeguards. In summary, model selection for psychiatric applications should weigh not only accuracy but also instruction-following ability, efficiency relative to size, potential domain adaptation, and the intended trade-off between safety and coverage.
Given the clinical importance and complexity of severity estimation, we conducted a targeted error analysis to better understand where models failed, focusing on symptom subtlety, contextual sensitivity, and diagnostic bias. Across models, the most common failure mode was under-recognition of mild or subthreshold symptom presentations; for example, depressive symptoms conveyed indirectly (“It’s been harder to get out of bed lately, but I’m fine”) were often misclassified as minimal severity even by high-performing models such as GPT-4o and Claude 3.5, highlighting the challenge of lexical subtlety and indirect phrasing. Contextual cues such as sarcasm, humor, or cultural idioms also led to misclassifications; in one illustrative case (“Guess I’m winning the ‘most anxious’ award this year”), models like Phi-3-mini interpreted the statement literally as mild anxiety rather than recognizing hyperbolic humor indicating moderate to severe symptoms, whereas contextually stronger models like GPT-4o and Qwen 2 72B were less affected but still occasionally erred. We also observed overpathologizing tendencies, particularly among smaller or less instruction-tuned architectures, where casual mentions of stress (“This week has been rough at work”) were classified as moderate depression or anxiety, potentially leading to unnecessary escalation. These patterns carry distinct clinical risks: missed mild symptoms may delay intervention, while overpathologizing can inflate false positives and burden services. Mitigating these errors may require context-aware pre-screening layers or hybrid human–AI review, especially for borderline cases.
Table 16: Error breakdown in disorder severity estimation. Frequencies represent the proportion of misclassified cases in the error analysis subset and sum to 100%.
Error Type Example Case Commonly Affected Models Frequency (%) Primary Risk in Deployment
Missed Mild Symptoms “It’s been harder to get out of bed lately, but I’m fine.” GPT-4o, Claude 3.5, Phi-3-mini 42 Under-intervention; delayed detection of early symptoms
Context Misinterpretation “Guess I’m winning the ‘most anxious’ award this year.” Phi-3-mini, Gemma 2 9B, Mistral NeMo 31 Misclassification; inaccurate triage in subtle cases
Overpathologizing “This week has been rough at work.” Phi-3-mini, Gemma 7B, smaller Llama 3 27 False positives; unnecessary escalation and resource strain
6.4 Nuanced Interpretation of Model Refusals
While refusal behavior in some models (e.g., Llama 3.1 family, Gemini) reduced the number of analyzable responses in our benchmarking, such behavior should not be interpreted solely as a limitation. In real-world psychiatric triage, refusal to answer may play a protective role preventing the generation of potentially harmful, misleading, or unsafe content.
We distinguish between undesirable refusals and desirable refusals:
  • Undesirable refusals occur when models decline to respond to benign evaluative prompts, such as classification of anonymized social media posts for research purposes. These refusals can hinder technical evaluation and limit model utility in controlled research settings.
  • Desirable refusals occur when the model detects that a task may produce unsafe or harmful outputs. For example, during an acute suicide risk scenario, a refusal to provide direct advice could allow the model to instead suggest immediate contact with emergency services or transfer to a human clinician. This aligns with clinical best practices, where safety and containment take precedence over diagnostic completeness.
From a deployment perspective, the optimal balance between responsiveness and safety will depend on context. Population-level screening tools may benefit from minimal refusals to ensure coverage, whereas crisis intervention systems may prioritize refusal in situations where any automated response could cause harm.
7 Limitations and Future Directions
Although this study offers valuable information on the capabilities of LLMs for mental health tasks, it is important to recognize several limitations that may affect the interpretation and generalizability of our findings. In addition to the future directions that can help achieve higher performance.
7.1 Limitations
These limitations arise from the inherent characteristics of the datasets we used, the constraints of the LLMs themselves, and the broader challenges associated with evaluating mental health through social media.
7.1.1 Dataset-Related Limitations
Data Source and Quality:
In this study, we relied on publicly available social media datasets to train and evaluate the LLMs. While these datasets offer a valuable resource for exploring mental health language, they lack the rigor of clinical reports (e.g., standardized diagnostic criteria). Though we prioritized datasets with human annotations [turcan2019dreaddit, mauriello2021sad, kabir2023deptweet], label subjectivity remains a concern – particularly for nuanced conditions like depression, where annotator agreement varies significantly [kabir2023deptweet]. Additionally, social media demographics over-represent younger, tech-literate populations, limiting generalizability to clinical populations. Our exclusive focus on English-language datasets further restricts cross-cultural applicability.
Limited Number of Disorders:
Our study focused on a limited set of mental health conditions, specifically stress, depression, and suicidal ideation, as detailed in Section 3.2. While these conditions are prevalent and important, they do not represent the full spectrum of mental health disorders. This limitation is primarily due to the availability of suitable datasets; datasets focusing on other mental health conditions that met our quality criteria were scarce. Consequently, our findings may not generalize to other mental health conditions, and further research is needed to evaluate the performance of LLMs across a broader range of disorders.
7.1.2 Model-Related Limitations
Context Length Restrictions:
Several models we evaluated, such as Phi-2, have limitations in the length of text they can process at once, constrained by their fixed context window sizes. While this was not a major limitation in our study as most datasets we used did not have particularly long text, it did prevent us from evaluating the CSSRS dataset. This limitation may pose challenges for future research involving longer social media posts or extensive patient histories.
API Filtering and Ethical Safeguards:
API filtering (e.g., Gemini 1.0 Pro blocking self-harm prompts) and built-in safeguards (e.g., Llama 3.1’s refusal rates) limited our ability to probe critical mental health scenarios. While ethically justified (to prevent misuse), these restrictions – exemplified by Llama 3.1’s 22% invalid response rate (Section 4.2.3) – create evaluation blind spots that may underestimate model risks in real-world deployment. This is the primary reason for excluding Gemini 1.5 for this study.
Model Stubbornness:
In our experiments, Some models (notably Mistral & Claude variants) frequently ignored formatting instructions (e.g., "answer concisely"), requiring iterative prompt adjustments (Section 3.5.2). This stubbornness complicated fair comparisons, as task performance metrics for these models may reflect compliance issues rather than true capability gaps.
7.1.3 Methodological Limitations
Focus on Specific Tasks:
Our study focused on a limited set of tasks: binary disorder detection, severity evaluation, and psychiatric knowledge assessment. While these tasks are important, they do not encompass the full spectrum of mental health applications. Our findings, therefore, may not be generalizable to other tasks, such as assessing empathy and engaging in therapeutic dialogue. Further research is needed to evaluate LLMs on a wider range of tasks to fully understand their capabilities and limitations in the context of mental healthcare.
Exploration of Prompt Space:
While we extensively explored various prompt engineering techniques, including role-playing and structured prompts, the vast space of possible prompts makes it impossible to guarantee that we identified the optimal ones for each model and task. Our results revealed that LLMperformance is highly sensitive to prompt wording and structure, with minor changes leading to significant result variations.
Computational Cost and Scalability:
Using large LLMs, such as GPT-4 and Claude 3 Opus, presented a high computational cost constraint in our study. To manage expenses, we employed sampling techniques, as discussed in Section 3.3, and relied on Hoeffding’s inequality (Equation 1) to determine appropriate sample sizes. While this allowed for a diverse range of models to be included, it limited the amount of data processed for each, meaning our findings are based on representative samples rather than complete datasets.
Lack of Transparency and Explainability:
The "black box" nature of LLMs limited our ability to investigate the reasoning behind model outputs. While this is a known methodological constraint of current architectures, it raises broader ethical concerns about deploying such systems in mental health contexts. Future research should focus on explainable AItechniques that can provide more transparent and interpretable results in the context of mental health evaluations.
@@ 7.2 Future Directions
Our analysis identifies six key priorities for advancing LLMapplications in mental health summarized in the following points:
  • First, improving model transparency is essential given the "black box" nature of current architectures. Implementing CoTprompting could mitigate this by enforcing explicit reasoning chains, given our finding that simple prompt changes caused 22.6% accuracy swings. This would help clinicians verify if models correctly weigh symptoms versus contextual factors in posts from platforms like Reddit.
  • Second, since smaller models (12B-14B parameters) matched or outperformed larger counterparts, future work should focus on integrating RAGwith clinical foundations. Linking models like Phi-3 and Mistral NeMo to diagnostic criteria from DSM-Vor anonymized hospital case notes (in closed clinical settings) would ground predictions in medical reality while enabling verifiable citations. This approach avoids costly parameter scaling while improving the clinical relevance.
  • Third, domain-specific fine-tuning using clinically vetted data is essential for accurate assessments. While our general-purpose models achieved up to 85% detection accuracy on certain datasets, their performance varied across datasets and disorders. Training open-source architectures on therapist-annotated case histories (with rigorous de-identification) could reduce these inconsistencies.
  • Fourth, expanding testing to non-English contexts across diverse demographics is crucial for clinical utility. While our work focused on English social media datasets, validating models on localized text (e.g., Arabic tweets or Korean forum posts) would assess cross-cultural robustness. Collaborations with multilingual healthcare providers are needed to create ethically sourced benchmarks without compromising patient anonymity.
  • Fifth, future research should assess multimodal LLMs and reasoning models such as o1, o3 and Deepseek-R1. Evaluating audio features (speech rhythm in depression) and video cues (facial affect recognition) alongside textual analysis may better approximate clinical evaluations.
    Finally, the field must expand beyond detection tasks to evaluate therapeutic uses. Systematic testing of LLMcapabilities in emotional support, empathy scoring, crisis de-escalation (e.g., suicide hotline simulations), and cognitive behavioural therapy style dialogue guidance will determine clinical readiness.
8 Conclusion
Our systematic evaluation of 33 LLMs across three clinically-relevant tasks—binary detection, severity estimation, and psychiatric knowledge assessment—reveals several key insights for mental health applications.
First, modern LLMs demonstrate strong binary classification capabilities, with GPT-4 achieving 85.2% accuracy on the SAD dataset and Mistral NeMo (12B parameters) outperforming larger architectures including GPT-4 (74.5% vs 72% on Dreaddit). However, reliability varies substantially across datasets and clinical conditions, with no single model family achieving universal superiority. Notably, the OpenAI ecosystem—particularly GPT-4—maintained the highest consistency, ranking as the top or near-top performer across all evaluation tasks.
Second, parameter count proves unreliable as a performance predictor. The 14B-parameter Phi-3-medium matched 80% of GPT-4’s cross-task capabilities, and Mistral NeMo dominated on the RED SAM dataset. Additionally, the 405B Llama 3.1 underperformed its older Llama 3 70B counterpart — a disparity we trace to excessive content filtering (22% valid query rejection rate) in the newer variant.
Third, methodological choices dramatically impact outcomes. Prompt engineering induced 22.6% accuracy variance in Gemma 7B, while three-example few-shot learning reduced Phi-3-mini’s severity estimation errors by 20%. These findings caution against comparing models without strict protocol standardization.
Fourth, in knowledge retrieval, more recently released models consistently outperformed predecessors, regardless of size, possibly due to architectural improvements or higher quality training data provided. The Llama 3.1 series dominated this category (405B: 91.2%, 70B: 89.5%), while Claude 3.5 Sonnet (82.5%) surpassed the older and larger Claude 3 Opus (81.8%). Even the modestly sized 14B Phi-3-medium achieved competitive performance (82.2%).
Three practical considerations emerge for implementing LLMs in mental healthcare: (1) Model architecture matters more than size, as shown by smaller models (e.g., Mistral NeMo) competing with much larger models – developers should prioritize task-specific evaluation over defaulting to the largest available models. (2) Standardized methods are crucial for reliable results, given that performance fluctuated significantly with different prompting techniques – the field needs consistent testing protocols to ensure valid model comparisons. (3) Ethical and practical needs should drive model choices – while closed-source systems like GPT-4 deliver strong benchmarks, open-source alternatives (e.g., Llama 3) offer needed transparency/customization despite balancing acts between safety controls and clinical effectiveness.
It is important to acknowledge the limitations of our study. While we focused on prevalent mental health conditions, our datasets, derived from social media, lack clinical validation and may reflect demographic biases. Evaluation constraints, such as APIfiltering and ethical guards (e.g., Llama 3.1’s query rejections) and model stubbornness (e.g., formatting issues with Claude and Mistral), could obscure real-world performance variations. Our English-only focus and computational sampling approach, driven by cost constraints, restrict the applicability of findings to diverse cultural contexts. Notably, the "black box" nature of LLMs complicates clinical interpretability despite their strong benchmarks.